|
Документ взят из кэша поисковой машины. Адрес
оригинального документа
: http://www.stsci.edu/smov/notes/0305smov.html
Дата изменения: Tue Mar 11 23:59:25 1997 Дата индексирования: Thu Feb 27 23:54:43 2014 Кодировка: Поисковые слова: http www.badastronomy.com phpbb index.php |
STATUS REPORTS
1. Observatory Status
MOSES status reports:
- PCS - There was guide star failure during the NICMOS 7134,
caused by the fact the select guide star turns out
to be a binary. This star is also used in the
subsequent NICMOS alignment activity (7041) in the
SMS069. It was agreed that the SMS will be rerun
with the K1 and K3 factors edited to produce the
effect of a lowered detection threshold, in an
effort to lock on to the star.
- All other observatory subsystems all nominal
2. WFPC2 Status
- Engineering status is nominal.
- Science status - nominal. The next UV throughput
measurements are scheduled for this evening,as are
the WFPC2 EROs and the resumption of Cycle 6 science.
3. NICMOS Status
- Engineering status - The instrument experienced another
SAA-induced transition to SUSPEND state about an hour
before the start of this meeting. The event, occurring
on Detector 3, has the same signature as the previous
anomalies. The memory dump, per procedure, is expected
shortly. Plans for proceding to the next NICMOS activity,
which is scheduled for March 11, will be determined after
more analysis of this event, and will be discussed at
this meeting tomorrow morning.
Yesterdays's Pre-alignment Check-out (7134) produced
nominal PAM motions, with the PAM motor temperature
peaking at around 30 degC. However, there were some
brief PRS limit violations caused by noise introduced
into the temperature monitors on the other motors. These
are to be expected; HSTARs were written.
- Science status - the Pre-Alignment Check-Out (7134) produced
the first NICMOS "engineering assessment frames." The
activity successfully verified correct postarg polarities
as well as the assumption, with only very small residual,
of 3.6 acrseconds of spacecraft motion per millimeter of PAM
defocus. The activity also verified assumptions regarding
exposure times and camera aperture locations (correct to
within 2 arcseconds).
The guide star failure, reported above, defaulted to one-
guide-star onctrol for the camera two data, but the
resulting loss of roll-control had no effect on the quality
of the data. The guide star failure resulted in no acqui-
sition at all for the Camera 3 observations, but, since the
drift produced smearing of only ~ 1/5 to 1/4 pixel, the data
are still valid. Therefore, the guide star problem had no
effect on the results of this activity, nor should it in the
upcoming 7041, in case of failure again there.
This activity produced images as the PAM swept throogh focus
in all three cameras. The Airy rings in cameras 1 & 2 are
as expected. Now the team will proceed with more
quantitative analysis, but as of now, they "like what they
see."
4. STIS Status
- Engineering Status - Nominal and unchanged - The MAMA LVPSs
remain off indefinitely, pending further investigation by
BASD. CCD radiator temperatures are running a few degrees
warmer now, possibly because of the increase in average sun
angle now that the BEA period is past.
- Science Status - Data from STIS internal observations
continue to be analyzed with updates expected for the
CCD bias values and aperture centers. The MSM positons
may not need updating.
Some SMOV activities will likely need to be repeated in some
way. These include some CCD flats, which were previously
saturated, the MSM tests (7081, 7083) which may be affected
by LVPS-induced temperatures, and a new proposal may be
required to test the thermal effects of a 12-hr on/off
cycling of the LVPS.
The first execution of the CCD UV monitor (7063) is not
usable because the pre-alignment offsets placed the star
at the edge of the 6x6 arcsec aperture. The plan for
subsequent visits in this proposal is to use a SIP on
March 9 to center the target in the aperture and to
provide a PDB SIAF update for the SMS072 in which the
next visit occurs.
5. FGS status - no new activity
6. FOC/COSTAR status- no new activity
PLANNING REPORTS
1. Status of Upcoming SMSs
- SMS062 - Currently executing.
- SMS065 - Signed off, ready for uplink.
- SMS069 - Final Mode products are out for review.
- SMS072 - Rebuilt without the NICMOS 7036 mentioned above.
- Final mode products are expected in ~ 1 day.
- SMS076 - Being reworked to include:
-- FGS ITS fix, which had failed
on the SMS062.
-- NICMOS 7135 modifications to avoid SAA orbits.
-- PAM motion spacings to reduce motor heating.
- SMS079 - Being reworked to include:
-- NICMOS 7054 SAA avoidance
2. Replans -None
OPERATIONS STATUS
1. OPUS
- Will prepare/support of unscheduled SIP on Mar. 9 to
recover STIS 7063 target mentioned above.
- Some NICMOS files went to trouble, last evening. Problem
is understood. In the meantime, data has been FTP's to IDT.
2. Archive
- For some observations, OPUS populated the association level
keywords for non-associated STIS data. This
confused the HDA and the REFERENCE file info for those
observations only was lost from the database tables. OPUS
has fixed the problem and we are working on fixing the
tables.